The shopkeeper brought a box containing x files for $20. he could have brought 5 more files if each file cost 20 cents less.(a) Form an equation in x and show that it reduces to x^2 + 5x - 500 = 0

1
Expert's answer
2021-11-08T13:22:07-0500

1$=100 cents

"\implies" 20 cents = 0.2 $

"\implies(\frac{20}{x}-0.2)(x+5)=20"

"(20-0.2x)(x+5)=20x"

"20x+100-0.2x^2-x=20x"

"0.2x^2+x-100=0"

"x^2+5x-500=0"